The number of typos in a book has a poisson distribution with an average of 5 typos
per 100 pages. What is the probability that you cannot find a typo in 50 pages.
1
Expert's answer
2020-04-20T10:22:33-0400

Poisson distribution,


"P(X=x)=\frac{\lambda^{x} e^{-\lambda}}{x!}"



"\lambda=5\ typos\ per\ 100 pages\\"

then it converts to 50 pages.

"\lambda=2.5\ typos\ per\ 50 pages"

The probability that you cannot find a typo in 50 pages"(P)=P(X=0)"

"P=\frac{2.5^{0} *e^{-2.5}}{0!}\\
\bold{P=0.0821}"